n small drops of same size are chargwd to V volts each .If they coalesce to form a single drop ,then its potential will be

Grade:11

1 Answers

Bhavit jain
7 years ago
V=kq/r Volume of small drop=4/3π r^3 Potential for big drop = k nq/R4/3π R^3 =n( 4/3 π r^3)R/(n)^(1/3 ) =r R=r (n)^(1/3 ) Potential for big drop = k nq/r (n)^(1/3 )